OpenStudy (anonymous):

hey guys what how do u answer this a+b+6=a+c

OpenStudy (anonymous):

ill give you a medal if u get it

OpenStudy (solomonzelman):

1) Are you sure the question says \(\large\color{black}{a+b+6=a+c }\) ? 2) What are you exactly asked to do?

OpenStudy (anonymous):

im helping someone on here

OpenStudy (anonymous):

There are three consecutive integers. The product of the smaller two integers added to 6 is equal to the product of the smallest and largest integer. What is the largest integer?

OpenStudy (anonymous):

thats the question

OpenStudy (solomonzelman):

`There are three consecutive integers.` \(\large\color{black}{(x-1)+x+(x+1) }\) (where x is the middle number.) ``` The product of the smaller two integers added to 6 is equal to the product of the smallest and largest integer. ``` \(\large\color{black}{(x-1)x+6=(x-1)(x+1) }\) Solve this equation.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x=7

OpenStudy (anonymous):

is that right

OpenStudy (solomonzelman):

``` (x-1)x+6=(x-1)(x+1) x^2-x+6=x^2-1 -x+6=-7 -x=-7 x=7 ```

OpenStudy (solomonzelman):

x is the middle number, so the largest number, the x+1, is 8.

OpenStudy (anonymous):

cool thanks

OpenStudy (solomonzelman):

Anytime!